OpenStudy (anonymous):

Find the remaining five trigonometric functions of θ. cotθ=(24/7), sinθ>0

OpenStudy (xapproachesinfinity):

what is cot? can you draw a picture of that

OpenStudy (xapproachesinfinity):

if cot=24/7 what tan would equal?

OpenStudy (xapproachesinfinity):

 this is your picture \(\rm tan\theta=7/24\)

OpenStudy (xapproachesinfinity):

find the rest

OpenStudy (anonymous):

Here is the problem.

OpenStudy (xapproachesinfinity):

no i'm talking about the picture that represents cot(theta)=24/7 i gave it to you? i gave you tan(theta) now find the rest

OpenStudy (xapproachesinfinity):

Hint use Pythagoras to find the remaining side so you will be able to find sin cos and then by that you will get sec and csc

OpenStudy (triciaal):